Air is approximately 21% oxygen. How many moles of oxygen are in 5.0 L of air at stp
1
Expert's answer
2019-09-16T04:05:45-0400

W% = V100%/Vtotal

V= Vtotal x W% / 100%

V= 5 x 21 /100 = 1.05 l

n= V/22.4 = 1.05 / 22.4 = 0.046875 moles